OpenStudy (hpfan101):

A rocket rises straight up from point X on the ground. A tracking device is on the ground 30 ft from point X. The distance from the tracking device to the rocket is changing at a rate of 100 ft/sec. At what rate is the height of the rocket changing when the rocket if 40 ft high?

OpenStudy (baru):



OpenStudy (baru):

you understand how i got that diagram?

OpenStudy (hpfan101):

Yeah I understand

OpenStudy (baru):

can you write 'd' in terms of 'h' ? hint: pythogoras theorem

OpenStudy (hpfan101):

\[d^2=h^2+30^2\] \[h=\sqrt{d^2-30^2}\]

OpenStudy (baru):

differentiate both sides with respect to 't'

OpenStudy (hpfan101):

\[\frac{ dh }{ dt }=\frac{ 1 }{ 2 }(d^2-30^2)^{-1/2}\times2d\]

OpenStudy (baru):

you have missed out a \(\frac{dd}{dt}\) term on the RHS

OpenStudy (hpfan101):

Oh okay.

OpenStudy (baru):

its easier to work with\[d^2=h^2+30^2\]

OpenStudy (baru):

can you differentiate that on both sides with respect to 't'?

OpenStudy (hpfan101):

Yeah \[2d \times \frac{ dd }{ dt }=2h \times \frac{ dh }{ dt }+0\]

OpenStudy (baru):

yep you are given that the rocket is 40 ft high, so h=40 can you find what 'd' is when h=40 (Pythagoras again)

OpenStudy (hpfan101):

Yeah d would be 50

OpenStudy (baru):

so now you have h=40 d=50 and dd/dt=100 (given in the question) substitute these values in the equation you just got after differentiating

OpenStudy (hpfan101):

\[2(50)\times100=2(40)\times \frac{ dh }{ dt }\] \[\frac{ dh }{ dt }=125ft/\sec \]

OpenStudy (baru):

yep, that's right :)

OpenStudy (hpfan101):

Cool! Thanks!